OpenStudy (anonymous):

Can someone explain to me how to do this equation? 10x^6+20x^3y^2/5x^3y

OpenStudy (anonymous):

what are you supposed to do with it?

OpenStudy (anonymous):

Im supposed to divide monomials

OpenStudy (anonymous):

maybe divide term by term

Parth (parthkohli):

orly\[{10x^6 + 20x^3y^2 \over 5x^3y} = {10x^6 \over 5x^3y} + {20x^3y^2 \over 5x^3y}\]

OpenStudy (anonymous):

PEMDAS- Please Exuce My Dear Aunt Sally (parenthases, exponents, multiplication, divition, audition, subtraction)

OpenStudy (anonymous):

\[\frac{10x^6}{5x^3y}=\frac{2x^{6-3}}{y}=\frac{2x^3}{y}\]

Parth (parthkohli):

Audition, subtraction? Well...

OpenStudy (anonymous):

ParthKohli. thank you. thats all i need to know :D

OpenStudy (anonymous):

banish that nonsense from your mind

Parth (parthkohli):

"Please excuse my dear aunt sally" is a longer way to memorize this stuff.

Parth (parthkohli):

Please don't memorize this like the above or your local cops will catch you.